# Efficient Data Collection for Robotic Manipulation via Compositional Generalization

### v1

@@ V-B Pairwise Composition
In this evaluation, we study pairwise composition of factors in BaseKitch. We compare L and No Variation to evaluate composition. For each pair of factors, we evaluate on all compositional combinations of factor values for that pair with respect to the L dataset, and set the values for other factors to their base values in $f^{*}$. This results in 9 evaluation scenarios for each of the 10 pairs. [^2] We have 3 values per factor that are not in the base values $f^{*}$, resulting in $93^{2}=$ 9 combinations for each of the (${5\choose 2}=10$ pairs. For L, we compare both training from scratch, and using BridgeData V2 as prior data. For No Variation, we only evaluate training with BridgeData V2.
@@
We report our results in Table I. We find that by training on L data and using BridgeData V2 as prior data, our policy is able to compose the factor values present in the L data to generalize to 59/90 of the compositional combinations of factor values. These results suggest that composition is possible in real-world robotic imitation learning, similar to our simulation results. However, unlike in simulation, leveraging prior data is critical for strengthening compositional abilities in real. Without prior data, the policy’s overall compositional success rate is roughly halved to 28/90, with success rates dropping for 9/10 pairs of factors. This resembles results in prior work that suggest end-to-end neural models can struggle with composition, but large pre-trained models can exhibit strong composition [51, 13]. We hypothesize that prior data may be needed in real, but not in simulation, due to additional minor factor variations inherent to real experiments, including for factors that we do not account for. Our greater emphasis on physical factors in real may also account for some of this discrepancy. However, prior data alone is not enough for generalization, as training on No Variation with prior data performs much worse than using L data, achieving a success rate of only 22/90. This suggests that with current prior robotic datasets, varied in-domain data is often still critical for generalization, motivating efficient in-domain data collection.
@@
We note that composition is generally strongest for pairs where at least one factor is visual, as opposed to physical factors that affect required physical motion. We hypothesize this is because physical factors can interact in more complex ways, making it more challenging for policies to compose unseen combinations of them. For example, object position and table height have the weakest composition together, possibly because both significantly affect the required grasp motion, and thus composition requires executing completely unseen grasps. Similarly, composition for container type is the weakest for the value white cup (second row in Fig. 7), which is narrower and taller than the other containers, and therefore requires a different place motion. In particular, this composes poorly with object position and table height, both physical factors. In contrast, composition is generally the strongest for object type. Although object type can sometimes affect required grasp and place motions, especially when interacting with other physical factors, we hypothesize it composes well because the different forks we consider are similar enough to not require drastically different physical motions. Overall, these results suggest that even with current prior robotics datasets, composition between physical factors can still be challenging, and as a result more exhaustive coverage for these factors during in-domain data collection may be necessary.
@@
Pre-trained Representations. Pre-trained visual representations have become popular for promoting generalization in robot learning. To assess their impact on composition, we additionally compare with using R3M, a representation intended for robotic manipulation, as a frozen visual encoder [33]. We evaluate on a subset of factor pairs from our complete pairwise evaluation. Specifically, we consider object type + table texture (the pair of factors from our complete evaluation that affect physical motion the least), object position + table texture (a pair where one factor is physical and the other is visual), and object position + table height (a pair where both factors are physical). In our results in Table II, we find that R3M performs poorly overall. Qualitatively, we observe that the trajectories are more jittery than when learning end-to-end, as also noted in prior work that tried R3M with diffusion-based policies [9]. This suggests that modern frozen visual representations are not effective for compositional generalization on real robots, and that end-to-end learning is superior for this.
@@
Data Strategy L No Variation Train Method Bridge R3M From Scratch Bridge R3M Object Type + Table Tex 8/9 1/9 4/9 2/9 0/9 Object Pos + Table Tex 5/9 0/9 1/9 1/9 0/9 Object Pos + Table Height 2/9 1/9 0/9 1/9 0/9 Overall 15/27 2/27 5/27 4/27 0/27
@@
TABLE II: Additional real robot pairwise composition results with comparison to R3M as a frozen visual representation.
@@
TABLE III: Out-of-domain transfer results to new kitchens CompKitch and TileKitch. We find that using both varied in-domain data from BaseKitch and BridgeData V2 as prior data is critical for effective transfer to these new kitchens. Stair outperforms L, although both achieve significant levels of transfer. Co-fine-tuning generally performs better than only fine-tuning.
@@ V-C Out-Of-Domain Transfer
To further demonstrate the utility of exploiting composition during data collection, we assess the transfer abilities of policies trained on our datasets to entirely new environments that capture some of the factor variety accounted for during data collection. We evaluate in two new kitchens, which we refer to as CompKitch and TileKitch. [^3] We name these kitchens based on the material of their countertops. See Fig. 13 in Section B-B for additional views that make this more apparent. These kitchens inherently have some completely out-of-distribution factor values from those studied in BaseKitch (e.g., table texture, shown in Fig. 8), including for some factors we do not account for during data collection (e.g., distractor objects, lighting). We first evaluate in these new kitchens with no further changes from the base factor values $f^{*}$ in BaseKitch, aside from their inherent shifts. In addition, to assess beyond pairwise composition for multiple factors, we also evaluate in these kitchens with additional combinations of factor shifts. We visualize these different conditions in Fig. 8. For each data collection strategy, we again compare using BridgeData V2 as prior data, and training from scratch. In addition to using prior data for both pre-training and co-fine-tuning (as done with the previously evaluated models), we also compare to only fine-tuning on in-domain data from a pre-trained BridgeData V2 policy, to assess the importance of co-fine-tuning for compositional generalization and transfer.
@@
We report our results in Table III. We find that BridgeData V2 is critical for transfer, as 2 of the 3 policies trained from scratch are unable to transfer at all, while the other only achieves a success rate of 11/40. When co-fine-tuning with BridgeData V2, having varied in-domain data is still needed for robust transfer, as No Variation only achieves a success rate of 1/40. These results indicate that this transfer setting represents a significant and challenging domain shift.
@@
Despite this, Stair and L both achieve significant levels of transfer, with success rates of 31/40 and 24/40, respectively. Stair generally outperforms L, as was the case in simulation. This suggests that policies that use prior data are able to effectively compose factor values seen in their training data to transfer to new settings that require composition. When fine-tuning from a pre-trained model, rather than co-fine-tuning, Stair and L produce policies that achieve some transfer, but less consistently than when co-fine-tuning, with reduced success rates of 21/40 and 15/40, respectively. In particular, the L policy fails to transfer to CompKitch, and both policies fail to transfer to Plastic Fork + Higher Table in both kitchens. This suggests that co-fine-tuning is generally a stronger method of leveraging prior data to facilitate composition than pre-training alone. Overall, we believe these results further suggest that policies can exhibit composition to generalize to unseen settings, our data collection strategies are sufficient to achieve some of this composition, and that using prior data is important for this composition to happen effectively.
@@ V-D Additional Experiments
Camera Position Composition. We do not consider camera position among the primary factors in our evaluation, due to the difficulty of varying this factor in a controlled manner. However, we have one additional third-person camera on our robot platform that we also collect data for, which allows us to separately study compositional generalization with this secondary camera. We visualize the difference in perspective with this secondary camera from our main camera in Fig. 10.
We focus on the composition of camera position with table texture. We evaluate in BaseKitch, but using the secondary camera for observations. We set table texture to the value white marble (third row in Fig. 7), as this was the value that was the most challenging for policies lacking variation for table texture to generalize to. We train a policy on the combination of two sub-datasets: our L dataset from the primary camera view, and the same L dataset but from the secondary camera view, with the exception of data for table texture. We compare against training on only one of these sub-datasets, which either lack data with the correct value for camera position, or the correct value for table texture, that is seen during evaluation. We again evaluate how using BridgeData V2 as prior data affects this composition. In our results in Table V, we find that yet again, policies that do not use prior data struggle to generalize. When using prior data, training on each sub-dataset individually results in a policy that sometimes generalizes, but not consistently. However, training on both sub-datasets together is able to achieve a perfect success rate, suggesting that the policy is able to effectively compose camera position with table texture.
@@ VI Discussion
Summary. We investigate the compositional abilities of visual imitation learning policies for robotic manipulation, and whether this can be exploited by data collection strategies to more efficiently achieve broad generalization. In summary, our simulated and real-world experiments suggest the following:
  • Robot policies do exhibit significant composition, although this does not always happen on a real robot.
@@
Limitations and Future Work. While we conduct thorough real-world experiments for evaluating compositional generalization, it would be interesting to investigate how these results scale to more complex robot platforms and tasks. We did some preliminary investigation into treating different tasks as a factor for composition, by combining data for a new task “remove fork from container” with the L dataset for our main task “put fork in container”. However, we were unable to achieve significant composition in early experiments. This could be because our new task data did not have any factor variation at all, which might be necessary for composition. Future work can more thoroughly consider this possibility.
@@
Although we believe our results can be broadly informative for data collection in robotics, we focus primarily on collecting in-domain data. Future work could scale this analysis to large-scale data collection efforts, to better understand how to most efficiently collect prior data for downstream transfer. While we do show that incorporating such prior data is critical for strengthening composition in our experiments, we only consider one prior dataset. Future work can investigate how other prior robotic datasets impact composition, and what aspects of prior datasets are the most important for this.
@@
Finally, while our results show that robot policies can possess significant composition and transfer capabilities, they still struggle at times with composition, particularly for factors that interact physically. Future work can better address generalization to these settings, such as through data collection strategies that emphasize certain factor combinations over others. Also, we only consider straightforward behavior cloning methods for policy learning. It would be interesting to study if other approaches to policy learning have better compositional abilities, such as different learning algorithms or architectures.
Appendix A Simulation Experiments
A-A Door Open Results
Figure 11: Simulation results of different data collection strategies for Door Open. We include results for every pair from 5 factors ($N=2$), average results across all pairs, and all 5 factors together ($N=5$). Results are similar as with Pick Place, with Stair generally performing the best, especially in the ($N=5$) setting. Error bars represent standard error across 5 seeds.
@@ A-A Door Open Results
We include additional simulation results for the task Door Open in Fig. 11. Results are similar to our results for Pick Place, with pairwise composition being strong for each pair of factors, and Stair generally performing the best, especially in the $N=5$ setting.
A-B Factors
For the inherently discrete-valued factors object texture, table texture, and distractor objects, we sample our $k$ discrete values from all possible training values specified in Factor World. Distractor objects also include a size scale (sampled from range $[0.3,0.8]$), 2D rotation (sampled from range $[0,2\pi]$), and 2D position (sampled from all possible positions on the table) as part of each value. For object position, we sample discrete 2D $xy$ positions from the range $[-0.1,0.1]$ for both coordinates. We note that the Pick Place task includes a small amount of added noise to object position each episode, sampled uniformly from the range $[-0.03,0.03]$. For camera position, we sample $xyz$ positions and rotation quaternions all from the range $[-0.05,0.05]$. When sampling our $k$ discrete values for each factor, we ensure that the scripted policy is able to solve the task for each value, because some values for object position and distractor objects can impede the task. In Fig. 12, we visualize two examples of values for each factor, for both the Pick Place and Door Open tasks. We note that each random seed in our evaluation includes a different set of $k$ discrete values sampled for each factor.
A-C Training
We use the same policy architecture and training hyperparameters used in the original Factor World experiments [46]. We condition policies on the same observations (2 84x84 RGB image observations from 2 different camera views without history, and proprioception). Unlike in the original Factor World experiments, we use random shift augmentations in all experiments. Unlike our real robot experiments, we train policies without goal conditioning, as we do not experiment with incorporating diverse prior data in this setting, which would benefit from some form of task conditioning.
A-D Computing Factor Changes
For each budget of total number of factor value changes, we determine the amount of factor value configurations allowed for each strategy. We then divide the budget of total demonstrations by this to determine how many demonstrations to collect for each configuration, with any remainder going to the last configuration.
@@ A-D Computing Factor Changes
Appendix B Real Robot Experiments
B-A Robot Platform
We use Logitech C920 webcams for our main and secondary third-person cameras. Our setup also includes the wrist camera used in BridgeData V2, and we experimented with using it in addition to our main third-person camera. While we found that using it improved overall robustness when training from scratch, particularly for some factors that wrist cameras provide invariance to (e.g., object position, table height), we also found that it was incompatible with using BridgeData V2 as prior data, reducing performance compared to not using prior data at all. We believe this is because only a small fraction of this dataset contains wrist camera data, as similar negative results with wrist cameras have been observed in other work that use prior robotic datasets with a small proportion of wrist camera data [34]. As we were only able to achieve successful transfer in our experiments in Section V-C by using prior data without a wrist camera, we decided to omit the wrist camera in all our experiments.

### v2

@@ V-B Pairwise Composition
In this evaluation, we study pairwise composition of factors in BaseKitch. We compare L and No Variation to evaluate composition. For each pair of factors, we evaluate on all compositional combinations of factor values for that pair with respect to the L dataset, and set the values for other factors to their base values in $f^{*}$. This results in 9 evaluation scenarios for each of the 10 pairs. [^2] We have 3 values per factor that are not in the base values $f^{*}$, resulting in $93^{2}=$ 9 combinations for each of the $\binom{5}{2}=10$ pairs. For L, we compare both training from scratch, and using BridgeData V2 as prior data. For No Variation, we only evaluate training with BridgeData V2.
@@
We report our results in Table I. We find that by training on L data and using BridgeData V2, our policy is able to generalize to 59/90 of the compositional factor value combinations. This suggests that composition is possible in real-world robotic imitation learning, similar to our simulation results. However, unlike in simulation, leveraging prior data is critical for strengthening composition. Without prior data, the policy’s overall compositional success rate is roughly halved to 28/90, with success rates dropping for 9/10 factor pairs. This resembles results in prior work that suggest end-to-end neural models can struggle with composition, but large pre-trained models can exhibit strong composition [54, 13]. We hypothesize that prior data may be needed in real, but not simulation, due to additional minor factor variations inherent to real experiments, including those for factors that we do not account for. Our greater emphasis on physical factors in real may also account for some of this discrepancy.
However, prior data alone is not enough for generalization, as training on No Variation with prior data performs much worse than using L data, achieving a success rate of only 22/90. This suggests that with current prior robotic datasets, varied in-domain data is often still critical for generalization, motivating the need for efficient in-domain data collection.
@@
We note that composition is generally strongest for pairs where at least one factor is visual. We hypothesize this is because physical factors can interact in more complex ways, making it more challenging for policies to compose unseen combinations of them. For example, object position and table height have the weakest composition together, possibly because both significantly affect the required grasp motion, and thus composition requires executing completely unseen grasps. Similarly, composition for container type is the weakest for the value white cup (second row in Fig. 7), which is narrower and taller than the other containers, and therefore requires a different place motion. In particular, this composes poorly with object position and table height, both physical factors.
In contrast, composition is generally the strongest for object type. Although object type may affect grasp and place motions, especially when interacting with other physical factors, we hypothesize it composes well because the different forks we consider are similar enough to not require drastically different motions. Overall, these results suggest that even with current prior robotics datasets, composition between physical factors can still be challenging, and thus more exhaustive coverage for these factors during data collection may be necessary.
We provide further pairwise composition results for additional factors object/container position and object orientation in Section C-D. In Section C-E, we provide additional analysis on how pairwise composition is affected by how similar new factor values are to the base factor values $f^{*}$.
@@
Pre-trained Representations. Pre-trained visual representations have become popular for promoting generalization in robot learning. To assess their impact on composition, we additionally evaluate using R3M [34] and VC-1 [32] as frozen visual encoders. We evaluate on a subset of factor pairs from our full pairwise evaluation. Specifically, we consider object type + table texture (the pair of factors that affect physical motion the least), object position + table texture (a pair where one factor is physical and the other is visual), and object position + table height (a pair where both factors are physical).
In our results in Table II, we find that both R3M and VC-1 perform poorly, with VC-1 failing to complete the task at all. Qualitatively, we observe that the trajectories are more jittery than when learning end-to-end, as also noted in prior work that tried R3M with diffusion-based policies [9]. This suggests that modern frozen visual representations are not effective for compositional generalization on real robots, and that end-to-end learning with prior robot data is superior for this.
@@
Data Strategy L No Variation Train Method Bridge R3M VC-1 From Scratch Bridge R3M VC-1 Object Type + Table Tex 8/9 1/9 0/9 4/9 2/9 0/9 0/9 Object Pos + Table Tex 5/9 0/9 0/9 1/9 1/9 0/9 0/9 Object Pos + Table Height 2/9 1/9 0/9 0/9 1/9 0/9 0/9 Overall 15/27 2/27 0/27 5/27 4/27 0/27 0/27
@@
TABLE II: Additional real robot pairwise composition results for our “put fork in container” task with R3M and VC-1.
@@
TABLE III: Out-of-domain transfer results to new kitchens CompKitch and TileKitch. We find that varied in-domain data from BaseKitch, and BridgeData V2 as prior data, are both critical for effective transfer to these new kitchens. Stair outperforms L, although both achieve significant levels of transfer. Co-fine-tuning generally performs better than only fine-tuning.
@@ V-C Out-Of-Domain Transfer
To further demonstrate the utility of exploiting composition during data collection, we assess the transfer abilities of policies trained on our datasets to entirely new environments that capture some of the factor variety accounted for during data collection. We evaluate in two new kitchens, which we refer to as CompKitch and TileKitch. [^3] We name these kitchens based on the material of their countertops. See Fig. 15 in Section C-B for additional views that make this more apparent. These kitchens inherently have some completely out-of-distribution factor values from those studied in BaseKitch (e.g., table texture, shown in Fig. 8), including for some factors we do not account for during data collection (e.g., distractor objects, lighting).
We first evaluate in these new kitchens with no further changes from the base factor values $f^{*}$ in BaseKitch, aside from their inherent shifts. In addition, to assess beyond pairwise composition for multiple factors, we also evaluate in these kitchens with additional combinations of factor shifts. We visualize these different conditions in Fig. 8. For each data collection strategy, we again compare using BridgeData V2 as prior data, and training from scratch. In addition to using prior data for both pre-training and co-fine-tuning (as done with the previously evaluated models), we also compare to fine-tuning on only in-domain data from a pre-trained BridgeData V2 policy, to assess the importance of co-fine-tuning for compositional generalization and transfer.
@@
We report our results in Table III. We find that BridgeData V2 is critical for transfer, as one of the two policies trained from scratch is unable to transfer at all, while the other only achieves a success rate of 11/40. When co-fine-tuning with BridgeData V2, having varied in-domain data is still needed for robust transfer, as No Variation only achieves a success rate of 1/40. These results indicate that this transfer setting represents a significant and challenging domain shift.
@@
Despite this, Stair and L both achieve significant levels of transfer, with success rates of 31/40 and 24/40, respectively. Stair generally outperforms L, as was the case in simulation. This suggests that policies that use prior data are able to effectively transfer to new settings that require composition. When fine-tuning from a pre-trained model, Stair and L produce policies that achieve some transfer, but less consistently than when co-fine-tuning, with reduced success rates of 21/40 and 15/40, respectively. In particular, the L policy fails to transfer to CompKitch, and both policies fail to transfer to Plastic Fork + Higher Table in both kitchens. This suggests that co-fine-tuning is generally superior for facilitating composition than pre-training alone. Overall, we believe these results further suggest that policies can exhibit composition to generalize to unseen settings, our data collection strategies are sufficient to achieve some of this composition, and that prior data is important for this composition to happen effectively.
@@ V-D Additional Experiments
Camera Position Composition. We do not consider camera position among the primary factors in our real evaluation, due to the difficulty of varying this factor in a controlled manner. However, we have one additional third-person camera on our robot platform that we also collect data for, which allows us to separately study compositional generalization with this secondary camera. We visualize the difference in perspective with this secondary camera from our main camera in Fig. 10.
We focus on the composition of camera position with table texture. We evaluate in BaseKitch, but using the secondary camera for observations. We set table texture to the value white marble (third row in Fig. 7), as this was the value that was the most challenging for policies lacking variation for table texture to generalize to. We train a policy on the combination of two sub-datasets: our L dataset from the primary camera view, and the same L dataset but from the secondary camera view, with the exception of data for table texture. We compare against training on only one of these sub-datasets, which either lack data with the value for camera position or table texture that is seen during evaluation. We again evaluate how BridgeData V2 affects this composition.
In our results in Table V, we find that yet again, policies that do not use prior data struggle to generalize. When using prior data, training on each sub-dataset individually results in a policy that sometimes generalizes, but not consistently. However, training on both sub-datasets together is able to achieve a perfect success rate, suggesting effective composition of camera position with table texture.
@@ VI Discussion
VI-A Summary
We investigate the compositional abilities of visual imitation learning policies for robotic manipulation, and whether this can be exploited by data collection strategies to more efficiently achieve broad generalization. In summary, our simulated and real-world experiments suggest the following:
  • Robot policies do exhibit significant composition, although the degree of composition varies across different factor pairs, with composition between physical factors generally being the most challenging.
@@ VI-B Limitations and Future Work
More Robot Platforms and Tasks. While we conduct thorough real-world experiments for evaluating compositional generalization, it would be interesting to scale our experiments and analysis to more complex robot platforms and tasks. We did some preliminary investigation into treating different tasks as a factor for composition, by training a policy on the combination of the L dataset for our main task “put fork in container”, and No Variation data for a new task “remove fork from container”. We then evaluated this on the new task with factor values found in the main task data, but it did not succeed. This could be because our new task data did not have any factor variation, which might be necessary for composition. Future work can more thoroughly consider this setting.
@@
Large-Scale Data Collection. Although we believe our results can be broadly informative for data collection in robotics, we focus primarily on in-domain data. Future work could scale this analysis to large-scale data collection efforts, to better understand how to most efficiently collect prior data for downstream transfer. While we do show that incorporating such prior data is critical for strengthening composition, we only consider one prior dataset. Future work can investigate how other prior robotic datasets impact composition, and what aspects of prior datasets are the most important for this.
@@
Improving Composition. While our results show that robot policies can possess significant composition and transfer capabilities, they still struggle at times with composition, particularly for factors that interact physically. Future work can better address generalization in these settings, such as by prioritizing data collection for certain factor combinations over others. Also, we only consider straightforward behavior cloning methods for policy learning. It would be interesting to study if other policy learning methods have better compositional abilities, such as different learning algorithms or architectures.
For example, one possible direction for future work could involve conditioning diffusion-based policies on individual environmental factors. Then, the score predictions for separate factors could be combined to produce a policy that more effectively achieves composition, similar to prior work in the context of text-to-image generation [29]. Prior work has studied combining score predictions in this manner via classifier-free guidance [18] in the context of goal-conditioned robotic policy learning [42], but this direction has not yet been studied for composing environmental factors.
Acknowledgments
This work was supported by DARPA W911NF2210214, ONR N00014-22-1-2293, NSF 1941722, Toyota Research Institute, and Other Transaction award HR00112490375 from the U.S. Defense Advanced Research Projects Agency (DARPA) Friction for Accountability in Conversational Transactions (FACT) program. We thank Suvir Mirchandani, Joey Hejna, and other members of the ILIAD lab for helpful discussions and feedback. We also thank Abraham Lee for advice on setting up our robot platform to work with BridgeData V2.
Appendix A Data Collection Strategies
A-A Pseudocode
We provide pseudocode for the data collection strategies proposed in Section III-C that are intended to exploit compositional generalization (Diagonal, L, Stair).
Algorithm 1 Diagonal
Input: scene S, factor values F (size $N$ factors $\times$ $k$ values)
for $j\leftarrow 1$ to $k$ do
     $f\leftarrow 0^{N}$
     for $i\leftarrow 1$ to $N$ do
         $f_{i}\leftarrow\textbf{F}_{ij}$
     end for
     SetFactors(S, $f$)
     CollectData(S)
end for
Algorithm 2 L
Input: scene S, factor values F (size $N$ factors $\times$ $k$ values), base factor values $f^{*}$ (size $N$ factors)
for $i\leftarrow 1$ to $N$ do
     $f\leftarrow f^{*}$
     for $j\leftarrow 1$ to $k$ do
         $f_{i}\leftarrow\textbf{F}_{ij}$
         SetFactors(S, $f$)
         CollectData(S)
     end for
end for
Algorithm 3 Stair
Input: scene S, factor values F (size $N$ factors $\times$ $k$ values), base factor values $f^{*}$ (size $N$ factors)
$f\leftarrow f^{*}$
for $j\leftarrow 1$ to $k$ do
     for $i\leftarrow 1$ to $N$ do
         $f_{i}\leftarrow\textbf{F}_{ij}$
         SetFactors(S, $f$)
         CollectData(S)
     end for
end for
Appendix B Simulation Experiments
B-A Door Open Results
@@
We include additional simulation results for the task Door Open in Fig. 11. Results are similar as in Pick Place, with generally strong pairwise composition, and Stair generally performing the best, especially in the $N=5$ setting.
B-B Factors
@@
For the inherently discrete-valued factors object texture, table texture, and distractor objects, we sample our $k=10$ values for $\mathcal{F}^{N}$ from all possible training values specified in Factor World. Distractor objects also include a size scale (sampled from range $[0.3,0.8]$), 2D rotation (sampled from range $[0,2\pi]$), and 2D position (sampled from all possible positions on the table) as part of each value. For object position, we sample 2D $xy$ positions from the range $[-0.1,0.1]$ for both coordinates. We note that the Pick Place task includes a small amount of added noise to object position each episode, sampled uniformly from the range $[-0.03,0.03]$. For camera position, we sample 3D $xyz$ positions and 4D rotation quaternions all from the range $[-0.05,0.05]$. When sampling our $k$ values for each factor, we ensure that the scripted policy is able to solve the task for each value, because some values for object position and distractor objects can impede the task. In Fig. 12, we visualize two examples of values for each factor, for both the Pick Place and Door Open tasks. We note that each random seed in our evaluation includes a different set of $k=10$ values sampled for each factor for $\mathcal{F}^{N}$.
B-C Training
We use the same policy architecture and training hyperparameters from the original Factor World experiments [49]. We condition policies on the same observations (2 84x84 RGB images from 2 camera views without history, and proprioception), and use the same action space (3D end-effector position deltas and open/close gripper). Unlike the original Factor World experiments, we always use random shift augmentation. Unlike our real robot experiments, we train policies without goal conditioning, as we do not leverage diverse prior data in this setting, so task conditioning is not as essential.
B-D Color Jitter Augmentation
Figure 13: Simulation results of data collection strategies for Pick Place, with additional color jitter augmentation. Overall performance across all strategies slightly improves, but the overall trends are similar to as without color jitter. Error bars represent standard error across 5 seeds.
We do not use color jitter augmentation in our main simulation experiments, because the original experiments for Factor World found this to reduce overall generalization, except for when training variation was very low [49]. However, here we include additional results with color jitter (in addition to random shift). We find that overall performance across all strategies does slightly improve, but the overall trends across strategies are similar to as without color jitter.
B-E Accounting for Factor Value Similarity
Figure 14: Performance of the Stair strategy for the Pick Place task, when composing object position and camera position. We find that selecting factor values using a similarity-aware strategy outperforms random selection (as done in our main simulation results).
In practice, data collection strategies should ideally account for similarity between factor values to improve generalization. For example, when training policies to be robust to the factor object type, it would be desirable to prioritize object diversity during data collection to generalize to as many objects as possible, rather than collecting data for overly similar objects. Here, we demonstrate how our proposed data collection framework could incorporate notions of factor similarity when available, through additional Pick Place experiments.
We consider the composition of object position and camera position in the $N=2$ setting, as these are the only factors we study in simulation where computing similarity/distance is straightforward (we use Euclidean distance for 3D positions, and angular distance for camera rotation quaternions). We modify Stair to choose factor values using a similarity-aware strategy, rather than randomly as before. We do this by running a $k$-medoids algorithm on the set of 10 values for each factor, to determine which $k$ values/medoids in the set minimize the sum of distances from each value to its nearest medoid. We determine $k$ according to what is permissible under different factor change budgets (e.g., for 10 total changes in the $N=2$ setting, each factor can have $k=\frac{10}{2}=5$ values).
In Fig. 14, we compare using $k$-medoids for factor value selection (orange), and random selection as done in our main results (gray). We evaluate using the same procedure as in our main simulation results in Section IV-B, where the aim is to generalize to all factor value combinations. We find that accounting for similarity does indeed improve upon random selection, although it does not match observing all factor values (green), which requires 20 factor changes.
We note that it often not straightforward to compute similarity metrics for other factors, and that while accounting for similarity can be easily incorporated into our data collection framework, it is mostly orthogonal to our study of composition. We believe it would be interesting for future work to investigate methods of computing similarity metrics for other factors, such as by using text and/or image embeddings, and leveraging this to further improve data collection.
B-F Computing Factor Changes
For each budget of factor value changes, we determine the amount of factor value configurations allowed for each strategy. We then divide the budget of total demonstrations by this to determine how many to collect for each configuration, with any remainder going to the last configuration.
Appendix C Real Robot Experiments